# Hadamard Encoded Row Column Ultrasonic Expansive Scanning (HERCULES) with Bias-Switchable Row-Column Arrays

### v1

@@ Abstract
Top-Orthogonal-to-Bottom-Electrode (TOBE) arrays, also known as bias-switchable row-column arrays (RCAs), allow for imaging techniques otherwise impossible for non-bias-switachable RCAs. Hadamard Encoded Row Column Ultrasonic Expansive Scanning (HERCULES) is a novel imaging technique that allows for expansive 3D scanning by transmitting plane or cylindrical wavefronts and receiving using Hadamard-Encoded-Read-Out (HERO) to perform beamforming on what is effectively a full 2D synthetic receive aperture. This allows imaging beyond the shadow of the aperture of the RCA array, potentially allows for whole organ imaging and 3D visualization of tissue morphology. It additionally enables view large volumes through limited windows. In this work we demonstrated with simulation that we are able to image at comparable resolution to existing RCA imaging methods at hundreds of frames per second. We validated these simulations by demonstrating an experimental implementation of HERCULES using a custom fabricated TOBE array, custom biasing electronics, and a research ultrasound system. Furthermore, we assess our imaging capabilities by imaging a commercial phantom, and comparing our results to those taken with traditional RCA imaging methods. Finally, we verified our ability to image real tissue by imaging a xenograft mouse model.
@@ 1 Introduction
In a previous work we demonstrated that we can use electrostrictive bias encoding to achieve transmit encoding and create an effective synthetic transmit aperture. This effective aperture allows us to make high resolution 2D images, that have the potential to even surpass the imaging quality of matrix probes [24]. We have also shown in previous work that we can perform volumetric photoacoustic imaging with RCAs by using selective biasing [27]. In this work we will build on these achievements by demonstrating that we can perform volumetric ultrasound imaging using bias-sensitive row-column arrays. Using bias encoding, we will create a virtual synthetic receive aperture reminiscent of a fully wired matrix probe. Furthermore, unlike traditional RCA imaging methods, this imaging scheme will be able to use cylindrically diverging waves to imaging beyond the shadow of the aperture of the probe. This imaging scheme is call Hadamard Encoded Row Column Ultrasonic Electronic Scanning (HERCULES).
2 Methods
2.1 Hadamard-Encoded Read-Out (HERO) Acquisition
@@ 2.2 Simulation
We hypothesize that HERCULES will provide imaging characteristics that are superior to traditional row-column array, specifically improvements in contrast and a reduction in artifacts. Additionally, we expect that we can achieve comparable SNR when imaging a point outside the shadow of the aperture as when imaging a point under the shadow if we use a diverging wave. To this end, we will compare 3 imaging schemes along with some variations in silico. These imaging schemes are: 1) RCA imaging using a walking virtual line source (VLS) for transmission [29]. 2) RCA imaging using tilted plane wave (TPW) transmissions [30]. 3) HERCULES imaging.
We use Field II [31, 32] to simulate point spread functions for each imaging scheme. We model our RCAs as a 2D array of physical elements and with and apply an apodization reflecting the bias states of the intersecting row & column. Receive signals are created by adding up the signals of every element for a given row or column. For a fair comparison we utilize the same array specifications for all simulations; these parameters are listed in Table 1. Additionally, we maintain a constant number of transmits for a consistent acquisition time and energy output.
@@ 2.3 Experiment
To further evaluate the different imaging schemes, we performed imaging ex-vivo in mice. Tumor-bearing mice were used for these experiments. Animal experiments were approved through the University of Alberta ACUC (AUP #3982 and #2994). SCID Hairless Outbred (SHO, Charles River) mice were inoculated with subcutaneous B16F10 mouse melanoma tumors in the hind flank. Once flank tumors had grown to a size of at least 5 mm in diameter, the hind section of the mice were then submerged in degassed water and imaged with our 8MHz TOBE arrays.
@@ 2.4 Processing
We use intensity modulated ultrasound receive data from the Vantage System, and manipulate it with the rest of our beamforming pipeline to reconstruct our images. The first step is to apply a matched filter that models our transmit waveform to allow us to recover the effective results of a single cycles transmit. We also apply the Hilbert transform to obtain the analytical signal, and improve the effects of compounding in later processing stages. For our HERCULES data we then multiply the inverse of our encoding Hadamard matrix across transmits, to recover acquisition data for every element of our 2D grid of receive elements. From this point we perform standard delay-and-sum beamforming[41] to reconstruct an intensity image. We generate low resolution images from every transmit and every receive element, and coherently compound these images together to obtain a final high resolution image. During experiments, we perform this beamforming at real-time imaging rates. We perform offline reconstruction with custom beamforming software and scripts to validate, and perform post-processing on our results. 2D images are normalized to their maximum values. 3D Volumes are rendered in 3D Slicer[42], with median filtering and thresholding.
To measure resolution we use the line targets present in the ATS 539 phantom. We cut an elevational slice from the imaged volume, and measure the lateral and axial resolution of the points in the image. Similarly, to measure contrast we use the anechoic cyst targets in the phantom. For our contrast metric we use generalized Contrast-to-Noise-Ratio (gCNR) [43] and compare histograms of the pixels inside and outside the cysts to measure their detectability. In this case we beamform rectangular volumes centered on the cyst, elongated in the y-direction, and extract concentric cylindrical regions that represent inside, and outside the cyst, without including the transition region.
@@ 4 Conclusion
In this work we were able to demonstrate a new imaging method that leveraged the properties of bias-sensitive row column arrays with bias encoding to perform high resolution volumetric structural imaging at real-time rates without mechanical scanning. This method has comparable if not superior imaging quality to other row-column imaging techniques, with the additional unique benefit of being able to image beyond the shadow of the aperture. These properties help to open the way for various clinical applications such as full organ imaging, and imaging volumes through limited apertures, such as cardiovascular imaging through the ribs, and intracranial imaging. Future work could include implementation of methods to reduce the required transmits to generate a volume. Volumetric flow, super-resolution, and elastography imaging could also build on this work.

### v2

@@ Abstract
Top-Orthogonal-to-Bottom-Electrode (TOBE) arrays, also known as bias-switchable row-column arrays (RCAs), allow for imaging techniques otherwise impossible for non-bias-switachable RCAs. Hadamard Encoded Row Column Ultrasonic Expansive Scanning (HERCULES) is a novel imaging technique that allows for expansive 3D scanning by transmitting plane or cylindrical wavefronts and receiving using Hadamard-Encoded-Read-Out (HERO) to perform beamforming on what is effectively a full 2D synthetic receive aperture. This allows imaging beyond the shadow of the aperture of the RCA array, potentially allows for whole organ imaging and 3D visualization of tissue morphology. It additionally enables view large volumes through limited windows. In this work we demonstrated with simulation that we are able to image at comparable resolution to existing RCA imaging methods at tens to hundreds of volumes per second. We validated these simulations by demonstrating an experimental implementation of HERCULES using a custom fabricated TOBE array, custom biasing electronics, and a research ultrasound system. Furthermore, we assess our imaging capabilities by imaging a commercial phantom, and comparing our results to those taken with traditional RCA imaging methods. Finally, we verified our ability to image real tissue by imaging a xenograft mouse model.
@@ 1 Introduction
In a previous work we demonstrated that we can use electrostrictive bias encoding to achieve transmit encoding and create an effective synthetic transmit aperture. This effective aperture allows us to make high resolution 2D images, that have the potential to even surpass the imaging quality of matrix probes in a 2D plane [24]. We have also shown in previous work that we can perform volumetric photoacoustic imaging with RCAs by using selective biasing [27]. In this work we will build on these achievements by demonstrating that we can perform volumetric ultrasound imaging using bias-sensitive row-column arrays. Using bias encoding, we will create a virtual synthetic receive aperture reminiscent of a fully wired matrix probe. Furthermore, unlike traditional RCA imaging methods, this imaging scheme will be able to use cylindrically diverging waves to imaging beyond the shadow of the aperture of the probe. This imaging scheme is call Hadamard Encoded Row Column Ultrasonic Electronic Scanning (HERCULES).
2 Methods
2.1 Hadamard-Encoded Read-Out (HERO) Acquisition
@@ 2.2 Simulation
We hypothesize that HERCULES will provide imaging characteristics that are comparable to traditional row-column array, with the potential to also result in a reduction in artifacts. Additionally, we expect that we can achieve comparable SNR when imaging a point outside the shadow of the aperture as when imaging a point under the shadow if we use a diverging wave. To this end, we will compare 3 imaging schemes along with some variations in silico. These imaging schemes are: 1) RCA imaging using a walking virtual line source (VLS) for transmission [29]. 2) RCA imaging using tilted plane wave (TPW) transmissions [30]. 3) HERCULES imaging.
We use Field II [31, 32] to simulate point spread functions for each imaging scheme. We model our RCAs as a 2D array of physical elements and apply an apodization reflecting the bias states of the intersecting row & column. Receive signals are created by adding up the signals of every element for a given row or column. For a fair comparison we utilize the same array specifications for all simulations; these parameters are listed in Table 1. Additionally, we maintain a constant number of transmits for a consistent acquisition time and energy output.
@@ 2.3 Experiment
To further evaluate the different imaging schemes, we performed imaging ex-vivo in mice. Tumor-bearing mice were used for these experiments. Animal experiments were approved through the University of Alberta ACUC (AUP #3982 and #2994). SCID Hairless Outbred (SHO, Charles River) mice were injected with subcutaneous B16F10 mouse melanoma tumors in the hind flank. Once flank tumors had grown to a size of at least 5 mm in diameter, the hind section of the mice were then submerged in degassed water and imaged with our 8MHz TOBE arrays.
@@ 2.4 Processing
We use intensity modulated ultrasound receive data from the Vantage System, and manipulate it with the rest of our beamforming pipeline to reconstruct our images. The first step is to apply a matched filter that models our transmit waveform to allow us to recover the effective results of a single cycles transmit. We also apply the Hilbert transform to obtain the analytical signal, and improve the effects of compounding in later processing stages. For our HERCULES data we then multiply the inverse of our encoding Hadamard matrix across transmits, to recover acquisition data for every element of our 2D grid of receive elements. From this point we perform standard delay-and-sum beamforming[41] to reconstruct an intensity image. We generate low resolution images from every transmit and every receive element, and coherently compound these images together to obtain a final high resolution image. During experiments, we perform this beamforming at real-time imaging rates. We perform offline volume reconstruction with custom beamforming software and scripts to validate, and perform post-processing on our results. 2D images are normalized to their maximum values. 3D Volumes are rendered in MATLAB [42], with smoothing and thresholding.
To measure resolution we use the line targets present in the ATS 539 phantom. We cut an elevational slice from the imaged volume, and measure the lateral and axial resolution of the points in the image. Similarly, to measure contrast we use the anechoic cyst targets in the phantom. For our contrast metric we use generalized Contrast-to-Noise-Ratio (gCNR) [43] and compare histograms of the pixels inside and outside the cysts to measure their detectability. In this case we beamform rectangular volumes centred on the cyst, elongated in the y-direction, and extract concentric cylindrical regions that represent inside, and outside the cyst, without including the transition region.
@@ 4 Conclusion
In this work we were able to demonstrate a new imaging method that leveraged the properties of bias-sensitive row column arrays with bias encoding to perform high resolution volumetric structural imaging at real-time rates without mechanical scanning. This method has comparable imaging quality to other row-column imaging techniques, with the additional unique benefit of being able to image beyond the shadow of the aperture. These properties help to open the way for various clinical applications such as full organ imaging, and imaging volumes through limited apertures, such as cardiovascular imaging through the ribs, and intracranial imaging. Future work could include implementation of methods to reduce the required transmits to generate a volume. Volumetric flow, super-resolution, and elastography imaging could also build on this work.